An isosceles triangle has two long sides and one short side. The short side is half the length of a long side. If the perimeter of the triangle is 15cm, find the length of the short side.
Answer by richwmiller(17219)

x+x+x/2=15
2.5x=15
x=6
x/2=3
6+6+3=15